# Batched Bayesian optimization with correlated candidate uncertainties

### v1

@@ 2.3.2 Additional strategies for numerical approximation of qPO acquisition scores
Techniques beyond Monte Carlo integration may also enable the estimation of qPO. If the posterior is a multivariate Gaussian, the acquisition score in Eq. 7 can be recast as an orthant probability through a change of variables, providing an alternative approach to predicting $\Pr(x^{*}=x_{i})$ that does not rely on posterior sampling (Azimi et al., 2010; Azimi, 2012). This orthant probability, defined in A.3, may be approximated analytically using a whitening transformation (Azimi, 2012). Additionally, Cunningham et al. (2013) propose an expectation propagation approach (Minka, 2001) to directly approximate orthant probabilities, and Gessner et al. (2020) introduce an integrator for truncated Gaussians that accurately estimates even small Gaussian probabilities. The orthant probabilities may also be estimated using numerical integration, which involves Cholesky decomposition followed by Monte-Carlo sampling (Genz, 1992). Alternative methods to estimate high-dimensional orthant probabilities (Miwa et al., 2003; Craig, 2008; Ridgway, 2016) may also be applied to the estimation of qPO acquisition scores. We consider the implementation of these methods future work.
@@ 3.1 Batched Bayesian optimization
A naive batch construction strategy is to select the top $b$ compounds based on an acquisition function designed for sequential BO. However, because the utility of a given selection depends on other selections, the top-$b$ approach does not guarantee an optimal batch in general (Garnett, 2023).
One general approach to batch design is to, as closely as possible, replicate the behavior of a sequential policy. Azimi et al. (2010) define an acquisition function that minimizes the discrepancy between the batch policy and sequential policy behavior. Batches may alternatively be constructed iteratively (myopically) by hallucinating the outcomes of previous selections in the batch. This enables the optimization of batch-level acquisition functions like multipoint expected improvement (Ginsbourger et al., 2010) and batch upper confidence bound (Desautels et al., 2014). However, this approach is typically limited to cases where the surrogate model can efficiently be updated with hallucinated or pending data. Further, batches that are constructed myopically only approximately optimize the batch-level acquisition function. In contrast, our approach does not rely on hallucination and can be applied to any posterior which can be sampled or modeled as a multivariate Gaussian.
Diversity heuristics have also been applied to prevent the selection of candidates that would provide minimal marginal information gain. Gonzalez et al. (2016) use local penalization to construct diverse batches. Determinantal point processes have also been used for batch diversification in discrete optimization (Kathuria et al., 2016; Nava et al., 2022). Nguyen et al. (2016) and Groves & Pyzer-Knapp (2018) model the objective landscape as mixture of Gaussians and acquire predicted local optima, aiding in the design of more diverse batches. Strategies based on diversity heuristics generally assume that diverse batches support exploration, not necessarily exploitation. We describe in Section 3.2 how our method does capture diversity, even with exploitation as the sole motivation.
Sequential acquisition functions that randomly select candidates by sampling, like Thompson sampling, can be extended to the batch case by increasing the number of random samples. Parallel Thompson sampling involves sampling from the model posterior and acquiring the optimum point from each sample (Hernández-Lobato et al., 2017). TS-RSR uses a similar methodology, but selects from each sample the point that minimizes a regret to uncertainty ratio (Ren & Li, 2024). Dai et al. (2022) extend neural Thompson sampling (Zhang et al., 2020) to the batch setting; each of $b$ randomly initialized neural networks informs the acquisition of one selection in the batch. The randomness inherent to Thompson sampling and related acquisition strategies is expected to contribute to the search’s exploration. qPO, in contrast, is exploitative in nature and intended to make selections deterministically.
@@ 3.2 Comparison with alternative batch strategies
We continue with an illustrative example to highlight how the proposed approach captures elements of diversity. Consider the following predictive distribution for $\mathbf{y}$ given $\mathbf{x}$:
@@
The probabilities of $x_{1}$, $x_{2}$, and $x_{3}$ being optimal are roughly 84%, 0%, and 16%, respectively. For $b=2$, our acquisition strategy would select $x_{1}$ and $x_{3}$, while a greedy strategy based only on the posterior mean vector would select $x_{1}$ and $x_{2}$. We might also expect a diversity-aware acquisition strategy to select $x_{1}$ and $x_{3}$ if we assume that the high covariance between $x_{1}$ and $x_{2}$ is due to design space similarity. This assumption is particularly valid for molecular applications when the surrogate model is a Gaussian process with a Tanimoto kernel. Because the Tanimoto kernel defines prior covariance based on structural similarity, structurally similar compounds will have higher covariance. Our method naturally captures this sense of diversity through model covariance without requiring clustering or other definitions of diversity that are not inherent to the model itself.
When Monte Carlo integration is used to approximate qPO, our method resembles Thompson sampling, particularly the parallel Thompson sampling (pTS) approach proposed by Hernández-Lobato et al. (2017). Both methods choose a batch of $b$ inputs by selecting the maxima of posterior samples.
Our approach differs from parallel Thompson sampling in two notable ways. First, in the case of $b=1$, Thompson sampling chooses candidates randomly with probability proportional to $\Pr(x^{*}=x_{i})$, allowing pTS to balance exploration and exploitation. qPO aims to choose candidates deterministically that maximize $\Pr(x^{*}=x_{i})$, making it a purely exploitative strategy. While forgoing exploration may sound deleterious in an iterative setting, empirically we observe that our approach outperforms pTS even in iterative settings (Section 4). Second, qPO generalizes to the $b\geq 2$ case differently from pTS. As described in Algorithm 2 of Hernández-Lobato et al. (2017), if the optimum input for a particular sample is already in the batch, then the second most optimal input is added (or the third if the second most optimal input is already in the batch too, and so forth). If the model makes highly correlated predictions, this will result in a batch filled with candidates that have highly correlated predictions. In the analytical example shown above in Eq. 12, pTS will be more likely to select $x_{1}$ and $x_{2}$ than $x_{1}$ and $x_{3}$. In contrast, once our method has added a first candidate to the batch, subsequent selections are chosen conditioned on existing selections not being optimal.
@@ 4.1 Baselines and evaluation metrics
We apply qPO to two model-guided searches of chemical libraries and compare its performance to common batch acquisition functions: pTS (Hernández-Lobato et al., 2017), upper confidence bound (UCB), greedy, and multipoint expected improvement (qEI) (Ginsbourger et al., 2010). We analyze two metrics—the fraction of the true top-$k$ acquired and the average oracle value of the acquired top-$k$—and observe that our method performs as well as or better than baselines according to both.
Both demonstrations use a Tanimoto Gaussian process surrogate model with a constant mean that operates on 2048-length count Morgan fingerprints. qPO is implemented following Eq. 11 using $M=10{,}000$. Pairs of candidates $\{x_{i},x_{j}\}$ for which $\Pr(x^{*}=x_{i})=\Pr(x^{*}=x_{j})$ are ranked by their predicted mean. This also applies to any case where $\Pr(x^{*}=x_{i})=0$ (see A.4.1).
The cost of sampling from a multivariate Gaussian posterior for $N$ candidates is $O(N^{3})$ due to the Cholesky decomposition (Vono et al., 2022). To alleviate the computational cost of sampling-based methods—qEI, pTS, and qPO—for large values of $N$, in each iteration we reduce the set of $N$ candidates to $10{,}000$ using a greedy metric. We then apply the respective acquisition strategy to select a batch from the smaller set of $10{,}000$ candidates. Because this may impact the exploration characteristics of pTS and qEI, we include an additional baseline (“random10k”) that randomly selects compounds from these top $10{,}000$ candidates. Experimental details can be found in A.4.
Figure 2: Optimization performance for the iterative discovery of antibiotic compounds. (A-C) Retrieval of the true top 0.5%, 1%, and 5%, representing 197, 393, and 1,962 top-performing compounds, respectively. (D-F) Average oracle value of the top 10, 50, and 100 acquired compounds, respectively, where a lower oracle value indicates greater antibiotic activity. Shaded regions denote $\pm$ one standard error of the mean across ten runs. Data is provided in A.5, Table 1. qPO, multipoint probability of optimality (ours); pTS, parallel Thompson sampling; qEI, multipoint expected improvement; UCB, upper confidence bound.
@@ 4.2 Application to antibiotic discovery
Our first demonstration applies Bayesian optimization to the retrospective identification of putative antibiotics with activity against Staphylococcus aureus. Wong et al. (2024) experimentally screened 39,312 compounds for growth inhibition of S. aureus. In this case study, we apply BO to retrospectively search this experimental dataset for compounds with the lowest reported mean growth of S. aureus, indicating greatest antibiotic activity. For each run, we randomly select an initial batch of 50 compounds and select 50 new compounds at each of 10 iterations.
@@
Figure 3: Batch diversity in the first iteration of a model-guided optimization loop for antibiotic discovery. Networks depict selected batches in Iteration 1, the first iteration after training on a randomly selected (random seed of 7) initial batch of 50 compounds with corresponding mean growth inhibition values from Wong et al. (2024). Nodes represent acquired compounds, and edges are drawn for any compound pair with Tanimoto similarity greater than 0.4. Nodes are positioned using the Fruchterman-Reingold force-directed algorithm (Fruchterman & Reingold, 1991). Histograms portray the distribution of Tanimoto similarity scores for all pairs in the selected batch. qPO and pTS most greatly promote diversity in this analysis, with UCB selecting the least diverse batch. qPO, multipoint probability of optimality (ours); pTS, parallel Thompson sampling; qEI, multipoint expected improvement; UCB, upper confidence bound.
Optimization performance of qPO and competitive baselines is shown in Figure 2. qPO performs best on average according to all metrics plotted in Figure 2. This demonstrates that, although our approach is motivated by pure exploration, it can empirically outperform those that intend to consider both exploitation and exploration.
@@
We also analyze batch diversity in this empirical setting to validate our hypothesis that qPO implicitly captures diversity (Section 3.2). The diversity of acquired batches across acquisition strategies is visualized in Figure 3 (details in A.4.4). qPO and pTS appear to obtain the most diverse selections, with UCB selecting the least diverse batch. While this visualization depicts acquired batches in a single iteration, these results indicate that qPO can achieve diversity without imposing randomness or diversity heuristics.

### v2

@@ 2.3.2 Additional strategies for numerical approximation of qPO acquisition scores
Techniques beyond Monte Carlo integration may also enable the estimation of qPO, for example, if one wants or needs to mitigate the computational cost of Monte Carlo sampling. Menet et al. 38 estimate the probability of optimality with almost-linear time complexity by assuming that Gaussian posterior predictions are uncorrelated. If the posterior is a multivariate Gaussian, the acquisition score in Eq. 6 can be recast as an orthant probability through a change of variables, providing an alternative approach to predicting $\Pr(x^{*}=x_{i})$ that does not rely on posterior sampling 39, 40. This orthant probability, defined in Section S3, may be approximated analytically using a whitening transformation 40. Additionally, Cunningham et al. 41 propose an expectation propagation approach 42 to directly approximate orthant probabilities, and Gessner et al. 43 introduce an integrator for truncated Gaussians that accurately estimates even small Gaussian probabilities. The orthant probabilities may also be estimated using numerical integration, which involves Cholesky decomposition followed by Monte-Carlo sampling 44. Alternative methods to estimate high-dimensional orthant probabilities 45, 46, 47 may also be applied to the estimation of qPO acquisition scores. We consider the implementation of these methods future work.
@@ 3.1 Batched Bayesian optimization
A naive batch construction strategy is to select the top $b$ compounds based on an acquisition function designed for sequential BO. However, because the utility of a given selection depends on other selections, the top-$b$ approach does not guarantee an optimal batch in general 11.
One general approach to batch design is to, as closely as possible, replicate the behavior of a sequential policy. Azimi et al. 39 define an acquisition function that minimizes the discrepancy between the batch policy and sequential policy behavior. Batches may alternatively be constructed iteratively (myopically) by hallucinating the outcomes of previous selections in the batch. This enables the optimization of batch-level acquisition functions like multipoint expected improvement 21 and batch upper confidence bound 22. However, this approach is typically limited to cases where the surrogate model can efficiently be updated with hallucinated or pending data. Further, batches that are constructed myopically only approximately optimize the batch-level acquisition function. In contrast, our approach does not rely on hallucination and can be applied to any posterior which can be sampled or modeled as a multivariate Gaussian.
Methods to improve diversity have also been applied to prevent the selection of candidates that would provide minimal marginal information gain. Gonzalez et al. 17 use local penalization to construct diverse batches. Determinantal point processes (DPPs) have also been used for batch diversification in discrete optimization 18; Nava et al. 27 demonstrate improved theoretical convergence rates by incorporating DPPs into parallel Thompson sampling. Nguyen et al. 19 and Groves and Pyzer-Knapp 20 model the objective landscape as mixture of Gaussians and acquire predicted local optima. Strategies based on diversity heuristics generally assume that diverse batches support exploration, not necessarily exploitation. We describe in Section 3.2 how our method does capture diversity, even with exploitation as the primary motivation.
Sequential acquisition functions that randomly select candidates by sampling, like Thompson sampling, can be extended to the batch case by increasing the number of random samples. Parallel Thompson sampling involves sampling from the model posterior and acquiring the optimum point from each sample 24. TS-RSR uses a similar methodology, but selects from each sample the point that minimizes a regret to uncertainty ratio 26. Dai et al. 25 extend neural Thompson sampling 48 to the batch setting; each of $b$ randomly initialized neural networks determines one selection in the batch. The randomness inherent to Thompson sampling and related strategies is expected to contribute to the search’s exploration. qPO, in contrast, is intended to make selections deterministically.
@@ 3.2 Comparison with alternative batch strategies
We continue with an illustrative example to highlight how the proposed approach captures diversity, a key contribution to qPO’s robustness. Consider the following predictive distribution for $\mathbf{y}$ given $\mathbf{x}$:
@@
The probabilities of $x_{1}$, $x_{2}$, and $x_{3}$ being optimal are roughly 84%, 0%, and 16%, respectively. For $b=2$, our acquisition strategy would select $x_{1}$ and $x_{3}$, while a greedy strategy based only on the posterior mean vector would select $x_{1}$ and $x_{2}$. A diversity-aware acquisition strategy would select $x_{1}$ and $x_{3}$ if we assume that the high covariance between $x_{1}$ and $x_{2}$ reflects design space similarity. This assumption is particularly valid for molecular applications when the surrogate model is a Gaussian process with a Tanimoto kernel. Because the Tanimoto kernel defines prior covariance based on structural similarity, structurally similar compounds will have higher covariance. Our method naturally captures this sense of diversity through model covariance without requiring clustering or other definitions of diversity that are not inherent to the model itself. However, qPO relies on the covariance as an indicator of design space similarity, and qPO may fail to capture diversity when applied to model posteriors with unreliable estimates of covariance.
When Monte Carlo integration is used to approximate qPO, our method resembles Thompson sampling, particularly the parallel Thompson sampling (pTS) approach proposed by Hernández-Lobato et al. 24 and Kandasamy et al. 28. Both methods choose a batch of $b$ inputs by selecting the maxima of posterior samples. Our approach differs from parallel Thompson sampling in two notable ways. First, in the case of $b=1$, Thompson sampling chooses candidates randomly with probability proportional to $\Pr(x^{*}=x_{i})$. qPO aims to choose candidates deterministically that maximize $\Pr(x^{*}=x_{i})$. However, because a closed-form solution for qPO acquisition scores does not exist, the implementation of our acquisition strategy is not purely deterministic. Second, qPO generalizes to the $b\geq 2$ case differently from pTS. As described in Algorithm 2 of Hernández-Lobato et al. 24, if the optimum input for a particular sample is already in the batch, then the second most optimal input is added (or the third if the second most optimal input is also in the batch, and so forth). If the model makes highly correlated predictions, this will result in a batch filled with candidates that have highly correlated predictions. In the analytical example shown above in Eq. 10, pTS will be more likely to select $x_{1}$ and $x_{2}$ than $x_{1}$ and $x_{3}$. In contrast, once our method has added a first candidate to the batch, subsequent selections are chosen conditioned on existing selections not being optimal.
@@ 4.1 Baselines and evaluation metrics
We apply qPO to two model-guided searches of chemical libraries and compare its performance to alternative batch acquisition functions: pTS 24, 28, Determinantal Point Process Thomspon Sampling (DPP-TS) 27, Thompson Sampling with Regret to Sigma Ratio (TS-RSR) 26, General-purpose Information-Based Bayesian OptimisatioN (GIBBON) 49, multipoint probability of improvement (qPI), multipoint expected improvement (qEI), batch upper confidence bound (BUCB) 50, upper confidence bound (UCB), and greedy (mean only). DPP-TS and TS-RSR are extensions of pTS with improved theoretical guarantees. BUCB, qPI, and qEI are batch-level extensions of sequential policies that employ myopic construction in finite discrete domains. GIBBON is an information-based batch acquisition function. We analyze the retrieval of the true top-$k$ acquired, the average oracle value of the acquired top-$k$, and computational cost (total run time), the former two being common metrics for assessing model-guided virtual screening methods 51, 13, 16. Cumulative regret values are reported in Section S4.
Both demonstrations use a Tanimoto Gaussian process surrogate model with a constant mean that operates on 2048-length count Morgan fingerprints. qPO is implemented following Eq. 9 using $M=10{,}000$. Pairs of candidates $\{x_{i},x_{j}\}$ for which $\alpha_{i}=\alpha_{j}$ are ranked by their predicted mean (6.1). The cost of sampling from a multivariate Gaussian for $N$ candidates is $O(N^{3})$ due to the Cholesky decomposition 33. To alleviate the computational cost of sampling-based methods—pTS, DPP-TS, TS-RSR, GIBBON, qPI, qEI, BUCB, and qPO—for large values of $N$, in each iteration we reduce the set of $N$ candidates to $10{,}000$ using a greedy metric. We then apply the respective strategy to select a batch from the smaller set of candidates. For qPO, this modifies the set of candidates x considered in Algorithm 1. Because this may impact the exploration characteristics of pTS and qEI, we include an additional baseline (“random10k”) that randomly selects compounds from these top $10{,}000$ candidates. Section S5 compares qPO performance when using this pre-filtering strategy with pre-filtering based on UCB. Experimental details can be found in 6.
@@ 4.2 Application to antibiotic discovery
Our first demonstration applies Bayesian optimization to the retrospective identification of putative antibiotics with activity against Staphylococcus aureus. Wong et al. 52 experimentally screened 39,312 compounds for growth inhibition of S. aureus. We search this dataset for compounds with the lowest reported mean growth of S. aureus, indicating greatest antibiotic activity. For each run, we randomly select an initial batch of 50 compounds and select 50 more in each of 10 iterations.
Optimization performance of qPO and baselines is shown in Table 1. While no single strategy outperforms all others in all metrics, qPO consistently performs on par with state-of-the-art baselines, with qPI appearing to be the most competitive baseline in this experiment. The retrieval of the top 0.5% is plotted for all iterations in Figure 3A for qPO and competitive baselines. We report the total run time for each method as an indication of computational cost, highlighting that qPO requires a similar computational cost to batch acquisition functions like pTS and TS-RSR but a greater cost compared to qPI in this demonstration.
Method Iteration Top 10 Average ($\downarrow$) Top 100 Average ($\downarrow$) Fraction Top 0.5% ($\uparrow$) Fraction Top 1% ($\uparrow$) Wall Time (s) ($\downarrow$)
BUCB 5 0.15 $\pm$ 0.02 0.51 $\pm$ 0.07 0.05 $\pm$ 0.01 0.06 $\pm$ 0.01 144 $\pm$ 1
DPP-TS 5 0.15 $\pm$ 0.00 0.71 $\pm$ 0.01 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00 1090 $\pm$ 38
GIBBON 5 0.18 $\pm$ 0.02 0.69 $\pm$ 0.04 0.02 $\pm$ 0.01 0.03 $\pm$ 0.01 392 $\pm$ 4
Greedy 5 0.18 $\pm$ 0.06 0.45 $\pm$ 0.07 0.05 $\pm$ 0.01 0.06 $\pm$ 0.01 103 $\pm$ 0
TS-RSR 5 0.18 $\pm$ 0.05 0.47 $\pm$ 0.08 0.05 $\pm$ 0.01 0.06 $\pm$ 0.01 0840 $\pm$ 10
UCB 5 0.19 $\pm$ 0.05 0.49 $\pm$ 0.08 0.05 $\pm$ 0.01 0.06 $\pm$ 0.01 103 $\pm$ 0
pTS 5 0.17 $\pm$ 0.04 0.62 $\pm$ 0.06 0.04 $\pm$ 0.01 0.05 $\pm$ 0.01 0865 $\pm$ 23
qEI 5 0.19 $\pm$ 0.04 0.62 $\pm$ 0.07 0.03 $\pm$ 0.01 0.04 $\pm$ 0.01 242 $\pm$ 3
qPI 5 0.14 $\pm$ 0.03 0.43 $\pm$ 0.07 0.06 $\pm$ 0.01 0.08 $\pm$ 0.01 145 $\pm$ 2
qPO 5 0.18 $\pm$ 0.06 0.46 $\pm$ 0.06 0.06 $\pm$ 0.01 0.07 $\pm$ 0.01 0859 $\pm$ 13
random10k 5 0.22 $\pm$ 0.02 0.80 $\pm$ 0.01 0.01 $\pm$ 0.00 0.01 $\pm$ 0.00 108 $\pm$ 6
BUCB 10 0.12 $\pm$ 0.02 0.27 $\pm$ 0.07 0.10 $\pm$ 0.02 0.13 $\pm$ 0.02 376 $\pm$ 1
DPP-TS 10 0.13 $\pm$ 0.00 0.55 $\pm$ 0.02 0.04 $\pm$ 0.00 0.04 $\pm$ 0.00 02687 $\pm$ 417
GIBBON 10 0.13 $\pm$ 0.01 0.51 $\pm$ 0.06 0.05 $\pm$ 0.01 0.05 $\pm$ 0.01 1096 $\pm$ 60
Greedy 10 0.11 $\pm$ 0.00 0.21 $\pm$ 0.02 0.11 $\pm$ 0.01 0.14 $\pm$ 0.02 105 $\pm$ 0
TS-RSR 10 0.11 $\pm$ 0.00 0.20 $\pm$ 0.03 0.12 $\pm$ 0.01 0.16 $\pm$ 0.02 1647 $\pm$ 16
UCB 10 0.11 $\pm$ 0.00 0.24 $\pm$ 0.05 0.12 $\pm$ 0.02 0.15 $\pm$ 0.02 105 $\pm$ 0
pTS 10 0.11 $\pm$ 0.01 0.29 $\pm$ 0.06 0.11 $\pm$ 0.02 0.13 $\pm$ 0.02 1696 $\pm$ 39
qEI 10 0.12 $\pm$ 0.01 0.32 $\pm$ 0.06 0.08 $\pm$ 0.01 0.11 $\pm$ 0.02 569 $\pm$ 3
qPI 10 0.10 $\pm$ 0.00 0.17 $\pm$ 0.01 0.14 $\pm$ 0.01 0.17 $\pm$ 0.01 375 $\pm$ 2
qPO 10 0.11 $\pm$ 0.01 0.21 $\pm$ 0.05 0.14 $\pm$ 0.02 0.19 $\pm$ 0.03 1684 $\pm$ 21
random10k 10 0.15 $\pm$ 0.00 0.68 $\pm$ 0.01 0.02 $\pm$ 0.00 0.02 $\pm$ 0.00 010 $\pm$ 6
Table 1: Optimization performance for the iterative discovery of antibiotic compounds at an intermediate and final iteration. The average oracle value of the top 10 and 100 acquired designs are shown, where a lower value indicates greater antibiotic activity. We also report retrieval of the true top 0.5% and 1%, representing 197 and 393 top-performing compounds, respectively, as well as wall time. All values denote averages $\pm$ one standard error of the mean across ten runs. The value(s) with the best mean performance by each metric is bolded, and remaining values whose mean are within one standard error of the mean of the best performer are underlined.
@@
Figure 2: Batch diversity of a model-guided optimization loop for antibiotic discovery. Networks depict selected batches in the first iteration after training on a randomly selected (seed of 7) initial batch of 50 designs with growth inhibition values from Wong et al. 52. Nodes represent acquired compounds; edges are drawn between pairs with Tanimoto similarity $>$ 0.4. Nodes are positioned using the Fruchterman-Reingold force-directed algorithm 53. Histograms portray the distribution of Tanimoto similarity scores for all pairs in the selected batch.
@@
We also analyze batch diversity to assess whether qPO captures diversity in this empirical setting. The diversity of acquired batches across select strategies is visualized in Figure 2 (details in 6.5). qPO, pTS, and qPI appear to obtain the most diverse selections, with UCB selecting the least diverse batch. While this visualization depicts acquired batches in a single iteration, these results indicate that qPO can achieve diversity without imposing randomness, myopic construction, or diversity heuristics.
Figure 3: Retrieval profile for two model-guided searches of chemical libraries. (A) Retrieval of the top 0.5% (197) designs for the iterative discovery of putative antibiotics (Section 4.2). (B) Retrieval of the top 0.01% (14) designs for the iterative discovery of organic materials (Section 4.3). For visibility, top-performing methods based on Tables 1 and 2 were selected for visualization. qPO performs on par with state-of-the-art methods for both case studies. qPI is competitive in both case studies, while TS-RSR is competitive primarily in the second case study (B). Shaded regions denote $\pm$ one standard error of the mean across ten runs.